Can you help me solve 2(x-8)=3(1-3x)-1/2
Answer by swincher4391(1107)

Distributive property is your friend here.
2(x-8) = 3(1-3x) - 1/2
The two on the left distributes to both terms, similarly the three on the right distributes to both terms.
2x - 2(8) = 3(1) - 3(3x) - 1/2
2x - 16 = 3 - 9x - 1/2
+9x + 9x
11x - 16 = 3 - 1/2
11x = 3 - 1/2 + 16
Multiply by 2 to get rid of the fraction.
22x = 6 - 1 + 32
22x = 37
x = 37/22